1.60 g O2 is reacted in excess S , if the percent yield of the reaction is 95.0% what is the yield of SO2
1
Expert's answer
2020-03-25T12:47:34-0400

The reaction between O2 and S can be shown as following:

S + O2 = SO2

The theoretical yield of SO2 produced in the reaction can be calculated using the equation:

m(SO2) = m(S) × Mr(SO2) / Mr(S) = 1.60g × 64.066 g/mol / 32.065 g/mol = 3.20 g

If the yield of the reaction is 95.0%, the actual yeald of SO2 is:

Actual yield = % yield × theoretical yield / 100% = 95.0% × 3.20 g / 100% = 3.04 g


Answer: 3.04 g